OpenStudy (anonymous):

I need an all around quickie on the gas laws- everything from Boyle's, Charles, Guy Lussac's, Molecular Kinetic Theory, ideal laws, etc... Any help would be greatly appreciated

OpenStudy (anonymous):

Ptotal = P1 + P2 + P3 ... Dalton’s Law of Partial Pressure X1 = n1/ntotal = P1/Ptotal Mole Fraction P1V1 = P2V2 Boyle’s Law V1 / T1 = V2 / T2 Charles’ Law V1 / n1 = V2 / n2 Avogadro's Law P1 / T1 = P2 / T2 Gay-Lussac’s Law P1V1 / T1 = P2V2 / T2 Combined Gas Law Graham's Law PV = nRT R = 8.3145 L kPa/mol K or R= 0.08206 L atm/mol K Ideal Gas Law (mm) P = dRT mm = molar mass d = density R= 0.08206 L atm/mol K Gas Density/Molar Mass vrms = √(3RT / M) M = molar mass in kg / mol R = 8.3145 J/mol K Root Mean Square Velocity

OpenStudy (anonymous):

i dont know very well whether that is what u really meant.

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (anonymous):

i hope it was of help to you